May is M years old now. Six years ago, April was four less than thrice May's age then. Find April's present age.
(Translating Word Phrases into Mathematical Expression)

Answer by mananth(16946)

May = m years
..
6 years ago
May = m-6 years
3 times m-6 = 3(m-6)
4 less
3(m-6)-4
...
..
April =3(m-6)-4
...
m.ananth@hotmail.ca